# Vision-Language Models Struggle to Align Entities across Modalities

### v1

@@ 7 Conclusions
Future work could focus on evaluating tasks with greater pointer or target attribute complexity. This benchmark can be easily extended to scenarios involving multiple pointer and target attributes (e.g., "Identify the name and rotation of the blue sphere") and/or a higher number of objects. The last alternative may be especially interesting to evaluate test-time scaling techniques, trying to better align with humans’ capability of using more time for more complex scenarios.
@@ Appendix B MATE Benchmark Dataset Format
Appendix C All Evaluated Model Results
@@
Appendix D Zero- and Few-shot Experiments
@@
Appendix E Linking Attribute Performance
@@
Appendix F Evaluation prompts
@@
F.1 UniModal Evaluation Prompts
@@
F.2 Cross-modal Evaluation Prompts
@@
F.3 Chain-of-Thought
The prompts used for the chain-of-thought experiments are shown in Figure 9 (for the image-to-text task) and 10 (for text-to-image).

### v2

@@ 7 Conclusions
Future work could focus on evaluating tasks with greater complexity regarding the pointer and target attributes. One of the main advantages of this benchmark is that it can be easily extended to scenarios involving multiple pointer and target attributes (e.g., "Identify the name and rotation of the blue sphere") and/or a higher number of objects. The last alternative may be especially interesting to evaluate test-time scaling techniques, trying to better align with humans’ capability of using more time for more complex scenarios.
Finally, it would be interesting to test models on cross-modal linking in a real-world scenario. One could choose a set of real-world entities (people, cars, etc.) and discretise a set of attributes that uniquely identify the entities (clothing, colour, size, etc). However, the main challenge for such an undertaking is to design the accompanying knowledge base such that you maintain a well-defined scope for the controlled evaluation of cross-modal entity linking.
@@ Acknowledgments
This work is partially supported by the Ministry of Science and Innovation of the Spanish Government (AWARE project TED2021-131617B-I00, DeepKnowledge project PID2021-127777OB-C21), project funded by MCIN/AEI/10.13039/501100011033 and by FEDER, the Basque Government (IXA excellence research group IT1570-22), the European Union under Horizon Europe (Project LUMINOUS, grant number 101135724), and the UK Engineering and Physical Sciences Research Council (grant EP/W002876/1).
@@ Appendix B MATE Benchmark Dataset Format
Appendix C Participants in Human Evaluation
The participants in the human evaluation ranged from 28 to 44 years old and all had university degrees.
Appendix D All Evaluated Model Results
@@
Appendix E Predicted Object Attribute Overlapping in 3D Coordinate-Only Linking Attribute Cases
This section analyzes cases where 3D coordinates are the only linking attribute that uniquely identifies the target object. In Figure 5, we report how many attributes of Qwen2.5-VL’s predicted objects match the target’s.
Figure 5: Number of matching attributes between Qwen2.5-VL’s predicted object and the target object in cases where 3D coordinates are the only uniquely identifying linking attribute. A value of ’4’ indicates that the model selected the correct object (i.e., all attributes matched), while ’–’ indicates predictions that could not be linked to any existing object.
Appendix F Zero- and Few-shot Experiments
@@
Appendix G Linking Attribute Performance
@@
Appendix H Evaluation prompts
@@
H.1 UniModal Evaluation Prompts
@@
H.2 Cross-modal Evaluation Prompts
@@
H.3 Chain-of-Thought
The prompts used for the chain-of-thought experiments are shown in Figure 10 (for the image-to-text task) and 11 (for text-to-image).